OpenStudy (anonymous):

Solve by the substition method 7x+3y=50 -6x+y=50

OpenStudy (saifoo.khan):

-4, 26

OpenStudy (cathyangs):

y=6x+50 7x+3(6x+50)=50 7x+18x+150=50 25x=-100 x=-4 7(-4)+3y=50 -28+3y=50 3y=78 y=26

OpenStudy (anonymous):

8,2

OpenStudy (saifoo.khan):

impossible.

OpenStudy (anonymous):

Why is that?

OpenStudy (saifoo.khan):

7x+3y=50 7(-4) + 3(26) = 50 Proved.

OpenStudy (saifoo.khan):

You must have changed the question or something by mistake.